# Denoising weak lensing mass maps with diffusion model: systematic comparison with generative adversarial network

### v1

@@ Abstract
Context. Weak gravitational lensing (WL) is the unique and powerful probe into the large-scale structures of the Universe. The ongoing and upcoming imaging surveys profit from the high image quality to accurately measure the distortions of distant galaxies and realise the unprecedented precision of the WL measurement. However, the finite number of source galaxies results in the dominant noise in the measurement, which is referred to as shape noise. Removing the shape noise from the observed WL field, i.e., denoising, enhances the potential of WL by accessing information at small scales where the shape noise dominates without denoising
Aims. Image-to-image translation methods, which are originally developed in the machine learning (ML) field, are a promising and suitable approach for denoising WL observations. We investigate the performance of denoising with ML-based approaches.
Methods. We utilise two ML models for denosing: generative adversarial network (GAN) and diffusion model (DM). Both methods have been successful in multiple tasks: inpainting, super-resolution, colourisation, etc. We evaluate the performance of denosing with GAN and DM utilising the large suite of mock WL observations, which serve as the training and test data sets.
Results. We apply denoising to 1,000 maps with GAN and DM models trained with 39,000 mock observations. Both models can fairly well reproduce the true convergence map on large scales. Then, we measure cosmological statistics: power spectrum, bispectrum, one-point probability distribution function, peak and minima counts, and scattering transform coefficients. We find that DM outperforms GAN in almost all statistics and recovers the correct statistics down to small scales within roughly $0.3\sigma$ level at the scales accessible from current and future WL surveys. We also conduct the stress tests on the trained model; denoising the maps with different characteristics, e.g., different source redshifts, from the data used in training. The performance degrades at small scales, but the statistics can still be recovered at large scales.
Conclusions. Though the training of DM is more computationally demanding compared with GAN, there are several advantages: numerically stable training, higher performance in the reconstruction of cosmological statistics, and sampling multiple realisations once the model is trained. It has been known that DM can generate higher-quality images in real-world problems than GAN, the superiority has been confirmed as well in the WL denoising problem.
Key Words.:
large-scale structure of Universe – Cosmology: observations – Methods: numerical
@@ 1 Introduction
Weak gravitational lensing (WL) is one of the major probes into the matter distribution in the Universe. Through the statistical analysis of the matter distribution mapped with WL, one can address fundamental physics such as the nature of dark matter and dark energy (for a review, see, e.g., Weinberg et al. 2013). Stage-III WL surveys, including Dark Energy Survey (DES; Dark Energy Survey Collaboration 2016), Subaru Hyper-Suprime Cam (HSC; Aihara et al. 2018; Dalal et al. 2023; Li et al. 2023), and Kilo Degree Survey (KiDS; Hildebrandt et al. 2017; Heymans et al. 2021; Wright et al. 2025), have already proven the capability of WL to constrain cosmological parameters precisely. Furthermore, Stage-IV WL surveys such as Euclid (Euclid Collaboration 2024), Vera C. Rubin Observatory Legacy Survey of Space and Time (LSST; LSST Science Collaboration 2009) and Nancy Grace Roman Space Telescope (Akeson et al. 2019) will enable the WL measurement with unprecedented precision with wide area coverage and deep imaging.
@@
WL refers to small distortions of distant galaxies, which are sensitive to the foreground gravitational potential. Through the measurement of the shapes of galaxies, one can reconstruct the projected density field. The analysis is referred to as mass reconstruction (Kaiser & Squires 1993; Schneider & Seitz 1995; Bartelmann 1995) and the resultant two-dimensional projected density field is called mass map. In mass reconstruction analysis, the number of source galaxies is finite, and thus, the intrinsic shape of source galaxies dilutes the WL signal. This noise is referred to as shape noise and is the major contaminant in mass maps. The statistical property of the shape noise is fairly known (van Waerbeke 2000), and various approaches to reduce the shape noise have been proposed (Bartelmann et al. 1996; Bridle et al. 1998; Marshall et al. 2002; Starck et al. 2006; Deriaz et al. 2012). Precise map-level reduction of shape noise, which is referred to as denoising, is critical for cluster search (Hamana et al. 2004; Shirasaki et al. 2015) and identification of filamentary structures (Higuchi et al. 2014).
@@
Recently, machine learning (ML) has emerged as a promising approach to accurate and fast mass reconstruction. In particular, generative models in the image domain, which take two-dimensional images as inputs and outputs, are suitable for denoising mass maps. For example, the generative adversarial network (GAN; Goodfellow et al. 2014) trained with mock WL simulations demonstrate high performance in denoising mass maps (Shirasaki et al. 2019) and surrogating gravitational evolution for WL (Yiu et al. 2022; Shirasaki & Ikeda 2024; Boruah et al. 2024). In this work, we focus on the diffusion model (DM; Sohl-Dickstein et al. 2015), which is one of the most powerful generative models. DM is also classified as a deep generative model and consists of two processes: the forward process, which adds noise to the input field, and the reverse process, which removes the noise from the noisy data. The forward process is analytically tractable, but the reverse process is not trivial, and deep neural networks are employed to model the reverse process. Starting from the white noise, the target data is reconstructed with iterations of the reverse process. DM can handle various tasks relevant to image processing, such as inpainting, super-resolution, and colourisation, and demonstrates higher performance compared with existing generative models (Dhariwal & Nichol 2021). In addition, DM has also been applied in mass reconstruction (Remy et al. 2023; Boruah et al. 2025).
We explore the capability of DM in denoising mass maps and conduct a systematic comparison with GAN. We employ the large suite of mock WL simulations $\kappa$TNG (Osato et al. 2021) to train the models. $\kappa$TNG data set consists of 40,000 mass maps and we manually add the shape noise to simulate the realistic WL observations. Finally, we construct 40,000 pairs of noiseless and noisy mass maps, and we use 39,000 maps for training and 1,000 maps for testing. In contrast to GAN, DM can sample multiple realisations of denoised maps from a single noisy mass map. For each noisy map, we sample five realisations with DM. In order to mimic the multiple sampling for GAN, we train GAN networks five times with different weight initialisations. Finally, there are 5,000 denoised maps both for GAN and DM. Then, we compare the denoised mass maps at a pixel level and address the effect on the peaks in the mass maps. Next, we measure the quantitative comparisons of cosmological statistics: angular power spectrum, bispectrum, one-point probability distribution function, peak and minima counts, and scattering transform coefficients. We also conduct stress tests, where denoising is applied to the modified test data, e.g., mass maps with source redshifts shifted, to evaluate the robustness and flexibility of the trained models.
@@ 2 Basics of weak lensing
Here, we review the basics of the formalism of gravitational lensing (for a review, see Bartelmann & Schneider 2001; Kilbinger 2015; Mandelbaum 2018). Images of distant galaxies are distorted due to the gravitational potential sourced from the intervening matter. This gravitational lensing effect by the large-scale structures of the Universe is referred to as cosmic shear. The distortion of galaxy shapes is characterised by the distortion matrix $A$:
@@
where $\bm{\theta}$ is the observed angular position of a source object, $\bm{\beta}$ is the true angular position, $\kappa$ is the convergence, $\gamma_{1}$ and $\gamma_{2}$ are the shear, and $\omega$ is the rotation. In the WL regime ($\phi/c^{2}\ll 1$, where $\phi$ is the gravitational potential), the convergence can be expressed as the projected density convolved with the distance kernel $W_{\kappa}(\chi)$ along the line-of-sight direction:
[…]
@@
$\tilde{\kappa}(\bm{\ell})=\frac{\ell^{2}_{1}-\ell^{2}_{2}+2i\ell_{1}\ell_{2}}{\ell_{1}^{2}+\ell_{2}^{2}}\tilde{\gamma}(\bm{\ell}),$ (6)
where $\bm{\ell}=(\ell_{1},\ell_{2})$ is the wave-vector in the angular space, and $\tilde{\kappa}$ and $\tilde{\gamma}$ are the convergence and shear in Fourier space. The convergence field can be obtained by applying the inverse Fourier transform to $\tilde{\gamma}$. In practice, the convergence field is directly estimated from galaxy shapes with Kaiser–Squires inversion (Kaiser & Squires 1993; Kaiser et al. 1995). Similarly to the shear field, the shape noise also affects the estimation of the convergence field. In the WL regime, the shape noise $N$ appears in the observed convergence field $\kappa_{\mathrm{obs}}$ in the additive form:
@@
where $\delta_{\mathrm{D}}^{2}$ is the two-dimensional Dirac delta function, $\sigma^{2}_{\varepsilon}$ is the variance of intrinsic shapes, and $n_{\mathrm{g}}$ is the number density of source galaxies. Regardless of the underlying probability distribution of the intrinsic shape, the mean of the shapes asymptotically follows Gaussian distribution according to the central limit theorem. The intrinsic variance is not tunable, and thus, the high source number density is critical to accurately estimate the mass map. However, conducting deep imaging surveys with higher source number density and wide areas is challenging due to the limited observing time. In this work, in order to realize accurate mass map reconstruction without demanding high source number density, we utilize the machine learning approaches to denoise the observed convergence map, i.e., remove the shape noise from the observed map.
@@ 3.1 Generative models for image-to-image translation
Our objective is to remove the shape noise from the observed convergence field. In other words, the goal is to find the optimal mapping of the noisy convergence field to the noiseless convergence field. To this end, we apply two different machine learning (ML) approaches: generative adversarial networks (GANs) and diffusion models (DMs). These approaches are classified as deep generative models, which learn the probability distribution function of the target observable given the training data sets. In particular, the denoising process can be regarded as the image-to-image translation problem, where the input image is noisy observed convergence maps and the output image is noiseless true convergence maps. The deep generative models are effective at image-to-image translation problems and demonstrate high performance in practical tasks such as image inpainting, colourisation, and segmentation. Recently, these methods have also been applied to the noise reduction task in astrophysics and cosmology (e.g., Shirasaki et al. 2019; Moriwaki et al. 2021; Flöss et al. 2024). Here, we present the basics of GANs and DMs.
@@ 3.1.1 Generative adversarial network
First, we review the algorithm of GAN (Goodfellow et al. 2014). In particular, our model learns the mapping conditioned on the input convergence map, and thus, we utilise the conditioned version of GAN, i.e., conditional GAN  [^1] In this work, we employ conditional GAN only and refer to conditional GAN as “GAN” for simplicity.. GANs consist of two competing networks: the generator $G$, which produces the fake images conditioned on the given image, and the discriminator $C$, which distinguishes fake and true images. At the same time, the discriminator compares the input image with the target image to determine whether the target image is a real image or a fake image generated by the generator. Through training with data sets, the model parameters $\theta$ are determined by optimising the Jensen–Shannon divergence and regularisation term:
@@
where $\mathbb{E}_{\bm{p}}$ denotes the averaging operation over the variable $\bm{p}$, $\bm{x}$ and $\bm{y}$ are the input and output data vector, respectively, and $\bm{z}$ is a random noise added to the lowest layer of the generator. Although it is possible to learn the transformation from the input to the output without noise $\bm{z}$, Gaussian noise $\bm{z}$ is added to prevent overfitting to the training data distribution and induce a diversity of output data. During training, gradient descent is applied alternately to optimise both $D$ and $G$ in each step. As described in the original implementation of GANs (Goodfellow et al. 2014), during training of $G$, instead of minimizing $\log(1-D(\bm{x},G(\bm{x},\bm{z})))$, we minimize $-\log D(\bm{x},G(\bm{x},\bm{z}))$. Furthermore, to slow down the learning rate of $D$, the loss function is divided by two during the optimisation of $D$. In summary, we alternately minimize the loss functions $L_{G}$ and $L_{D}$, which are defined as
@@
In this work, we employ pix2pix (Isola et al. 2017), which implements GAN with PyTorch (Paszke et al. 2019) to take two-dimensional images as data vectors. The details of the architectures of the networks are described in Appendix A.1.
While GANs can generate high-quality images quickly, they suffer from various issues due to the model structure. In general, the training of GANs is unstable due to the mode collapse and the vanishing gradient problem compared with other existing generative models such as variational auto-encoders (Kingma & Welling 2014). Furthermore, the generated images are less diverse because the generator tends to ignore the noise term and generate images only from the information of the conditioned images as training proceeds. Due to this problem, pix2pix does not provide the noise term in the input data vectors. Instead, the stochasticity appears only in dropouts of the generator network. However, the output images are less diverse even with the treatment than other generative models, that is, GANs fail to learn the probability distribution of the target images.
@@ 3.1.2 Diffusion model
DM (Sohl-Dickstein et al. 2015; Ho et al. 2020; Song et al. 2021) has been proposed in order to circumvent the problems of GANs and enhance the quality of generated data. It is known that training of DM is more stable and can generate more diverse outputs from a single input data.
@@
where $\mathcal{N}(\bm{\mu},\bm{\Sigma})$ is the normal distribution with the mean $\bm{\mu}$ and the variance $\bm{\Sigma}$, and $\bm{I}$ is the identity matrix. The probability distribution function $q(\bm{y}_{t}|\bm{y}_{t-1})$ of the noise-added data $\bm{y}_{t}$ given the data at the previous step $\bm{y}_{t-1}$ is expressed as
@@
where the mean $\bm{\mu}_{\theta}$ and the variance $\bm{\Sigma}_{\theta}$ are expressed as the deep neural networks with model parameters $\theta$. Then, let us consider the consistency of the joint probability distribution function of the series of the data $(\bm{y}_{1},\cdots,\bm{y}_{T})$. The joint probability distributions are given with forward process or reverse process:
$q(\bm{y}_{1},\cdots,\bm{y}_{T})$ $=q(\bm{y}_{0})\prod_{t=1}^{T}q(\bm{y}_{t}|\bm{y}_{t-1}),$ (19)
$q_{\theta}(\bm{y}_{1},\cdots,\bm{y}_{T})$ $=p(\bm{y}_{T})\prod_{t=1}^{T}q_{\theta}(\bm{y}_{t-1}|\bm{y}_{t}),$ (20)
where $p(\bm{y}_{T})$ is the prior of the final data and taken as unit normal distribution. The objective is to minimise the Kullback–Leibler divergence of the two distributions: $q(\bm{y}_{1},\cdots,\bm{y}_{T})$ and $q_{\theta}(\bm{y}_{1},\cdots,\bm{y}_{T})$.
In practice, for better convergence of the training process and sampling efficiency, Ho et al. (2020) proposed the simplified scheme; the objective function for optimisation is the reweighted variational lower-bound and the variance is diagonal $\bm{\Sigma}_{\theta}=\sigma_{t}^{2}\bm{I}$. Finally, the loss function $L_{t}$ for the step $t$ is given as
@@
In this work, we employ Palette (Saharia et al. 2021) model, which implements DM to take images as inputs and outputs with PyTorch (Paszke et al. 2019). Similarly to GANs, Palette takes the input image $\bm{x}$ used as the conditioned image, and thus, the model should be referred to as “conditional” DM. Conditional DM can be implemented by adding the input image as the argument of the noise vector, i.e.,
@@
DM can generate more high-quality and diverse data compared with GAN. Furthermore, the training process of DM is more stable because of the simple objective function (Eq. 21). The drawback of DM is the computational cost of data generation. In general, the number of total diffusion steps $T$ is required to be $T\simeq 10^{3}$, which takes considerably longer than GANs.

### v2

# gravitational lensing: weak — large-scale structure of universe — cosmology: observations — methods: numerical

@@ Abstract
Weak gravitational lensing (WL) is the unique and powerful probe into the large-scale structures of the Universe. Removing the shape noise from the observed WL field, i.e., denoising, enhances the potential of WL by accessing information at small scales where the shape noise dominates without denoising. We utilise two machine learning (ML) models for denosing: generative adversarial network (GAN) and diffusion model (DM). We evaluate the performance of denosing with GAN and DM utilising the large suite of mock WL observations, which serve as the training and test data sets. We apply denoising to 1,000 noisy mass maps with GAN and DM models trained with 39,000 mock observations. Both models can fairly well reproduce the true convergence map on large scales. Then, we measure cosmological statistics: power spectrum, bispectrum, one-point probability distribution function, peak and minima counts, and scattering transform coefficients. We find that DM outperforms GAN in almost all considered statistics and recovers the correct statistics down to small scales. For example, the angular power spectrum can be recovered with DM up to multipoles $\ell\lesssim 6000$ while the noise power spectrum dominates from $\ell\simeq 2000$. We also conduct stress tests on the trained model; denoising the maps with different characteristics, e.g., different source redshifts, from the training data. The performance degrades at small scales, but the statistics can still be recovered at large scales. Though the training of DM is more computationally demanding compared with GAN, there are several advantages: numerically stable training, higher performance in the reconstruction of cosmological statistics, and sampling multiple realisations once the model is trained. It has been known that DM can generate higher-quality images in real-world problems than GAN, the superiority has been confirmed as well in the WL denoising problem.
@@ 1 Introduction
Weak gravitational lensing (WL) is one of the major probes into the matter distribution in the Universe. Through the statistical analysis of the matter distribution mapped with WL, one can address fundamental physics such as the nature of dark matter and dark energy (for a review, see, e.g., Weinberg2013). Stage-III WL surveys, including Dark Energy Survey (DES; DES2016), Subaru Hyper-Suprime Cam (HSC; Aihara2018; Dalal2023; Li2023), and Kilo Degree Survey (KiDS; Hildebrandt2017; Heymans2021; Wright2025), have already proven the capability of WL to constrain cosmological parameters precisely. Furthermore, Stage-IV WL surveys such as Euclid (Euclid), Vera C. Rubin Observatory Legacy Survey of Space and Time (LSST; LSST) and Nancy Grace Roman Space Telescope (WFIRST) will enable the WL measurement with unprecedented precision due to wider area coverage and deeper imaging.
@@
WL refers to small distortions of shapes of distant galaxies, which are sensitive to the foreground gravitational potential. Through the measurement of the galaxy shapes, one can reconstruct the projected density field. The analysis is referred to as mass reconstruction (Kaiser1993; Schneider1995; Bartelmann1995) and the resultant two-dimensional projected density field is called mass map. In mass reconstruction analysis, the number of source galaxies is finite, and thus, the intrinsic shape of source galaxies dilutes the WL signal due to the Poissonian nature. This noise is referred to as shape noise and is the major contaminant in mass maps. The statistical property of the shape noise is well studied (e.g., vanWaerbeke2000), and various approaches to reduce the shape noise have been proposed (Bartelmann1996; Bridle1998; Marshall2002; Starck2006; Deriaz2012). Precise map-level reduction of shape noise, which is referred to as denoising, is critical for cluster search (Hamana2004; Shirasaki2015) and identification of filamentary structures (Higuchi2014).
@@
Recently, machine learning (ML) has emerged as a promising approach to accurate and fast mass reconstruction (e.g., Jeffrey2020). In particular, generative models in the image domain, which take two-dimensional images as inputs and outputs, are suitable for denoising mass maps. For example, the generative adversarial network (GAN; Goodfellow2014) trained with mock WL simulations demonstrate high performance in denoising mass maps (Shirasaki2019; Whitney2024) and surrogating gravitational evolution for WL (Yiu2022; Shirasaki2024; Boruah2024). In this work, we focus on the diffusion model (DM; Sohl-Dickstein2015), which is one of the most powerful generative models. DM is also classified as a deep generative model and consists of two processes: the forward process, which adds noise to the input field, and the reverse process, which removes the noise from the noisy data. The forward process is analytically tractable, but the reverse process is not trivial, and deep neural networks are employed to model the reverse process. Starting from the white noise, the target data is reconstructed with iterations of the reverse process. DM can handle various tasks relevant to image recognition, such as inpainting, super-resolution, and colourisation, and demonstrates higher performance compared with existing generative models (Dhariwal2021). In addition, DM has also been applied in mass reconstruction (Remy2023; Boruah2025).
[…]
@@ 2 Basics of weak lensing
Here, we review the basics of the formalism of gravitational lensing (for a review, see Bartelmann2001; Kilbinger2015; Mandelbaum2018). Images of distant galaxies are distorted due to the gravitational potential sourced from the intervening matter. This gravitational lensing effect by the large-scale structures of the Universe is referred to as cosmic shear. The distortion of galaxy shapes is characterised by the distortion matrix $A$:
@@
where $\bm{\theta}$ is the observed angular position of a source object, $\bm{\beta}$ is the true angular position, $\kappa$ is the convergence, $\gamma_{1}$ and $\gamma_{2}$ are the shear, and $\omega$ is the rotation. In the WL regime ($\phi/c^{2}\ll 1$, where $\phi$ is the gravitational potential and $c$ is the speed of light), the convergence can be expressed as the projected density convolved with the distance kernel $W_{\kappa}(\chi)$ along the line-of-sight direction:
$\kappa(\bm{\theta})$ $=\int^{\chi_{\mathrm{H}}}_{0}d\chi\,W_{\kappa}(\chi)\,\delta_{\mathrm{m}}(r(\chi)\bm{\theta},\chi),$ (2)
$W_{\kappa}(\chi)$ $=\frac{3\Omega_{\mathrm{m}}H^{2}_{0}}{2c^{2}}r(\chi)\int^{\chi_{\mathrm{H}}}_{\chi}d\chi^{\prime}\,p(\chi^{\prime})\frac{r(\chi^{\prime}-\chi)}{r(\chi^{\prime})},$ (3)
where $\chi$ is the comoving distance, $\chi_{\mathrm{H}}$ is the comoving distance to the horizon, $\Omega_{\mathrm{m}}$ is the matter density at the present Universe, $H_{0}$ is the Hubble constant, $r(\chi)$ is the angular diameter distance, and $p(\chi)$ is the source galaxy distribution, which is normalized as $\int d\chi\,p(\chi)=1$. The convergence field is the tracer of the cosmic matter density field, and thus, the convergence map is also called a “mass map”. At the peak positions of high convergence, massive structures such as galaxy clusters are likely to be located. Hence, the peaks of the convergence map are used to search massive clusters (e.g., Miyazaki2018; Oguri2021; Chen2025).
@@
where $\langle\cdots\rangle$ denotes the ensemble average. To estimate the shear field in observations, we take averages of close galaxies which are gravitationally lensed with the same shear field. Since the number of close galaxies is finite, the average of the intrinsic shape is not necessarily zero and the residual dilutes the WL signal. This residual noise is called as “shape noise” and is the primary contaminant in WL measurements.
@@
where $\bm{\ell}=(\ell_{1},\ell_{2})$ is the wave-vector in the angular space, and $\tilde{\kappa}$ and $\tilde{\gamma}$ are the convergence and shear in Fourier space. The convergence field can be obtained by applying the inverse Fourier transform to $\tilde{\gamma}$. In practice, the convergence field is directly estimated from galaxy shapes with Kaiser–Squires (KS) inversion (Kaiser1993; Kaiser1995). Similarly to the shear field, the shape noise also affects the estimation of the convergence field. In the WL regime, the shape noise $N$ appears in the observed convergence field $\kappa_{\mathrm{obs}}$ in the additive form:
@@
where $\delta_{\mathrm{D}}^{2}$ is the two-dimensional Dirac delta function, $\sigma^{2}_{\varepsilon}$ is the variance of intrinsic shapes, and $n_{\mathrm{g}}$ is the number density of source galaxies. Regardless of the underlying probability distribution of the intrinsic shape, the mean of the shapes asymptotically follows Gaussian distribution according to the central limit theorem. The intrinsic variance is not tunable, and thus, the high source number density is critical to accurately estimate the mass map. However, conducting deep imaging surveys with higher source number density and wide areas is challenging due to the limited observing time. In this work, in order to realise accurate mass map reconstruction without demanding a high source number density, we utilise the machine learning approaches to denoise the observed convergence map, i.e., remove the shape noise from the observed map. There is a caveat about the statistical property of the shape noise. In the real WL measurements, it is possible to estimate the shape noise by randomly rotating source galaxies, which reflects the spatial variations of the number density of source galaxies. However, the assumed uniform noise does not depend on any specific source galaxy distribution except the number density, and thus, the results of denoising WL maps are more generic.
@@ 3.1 Generative models for image-to-image translation
Our objective is to remove the shape noise from the observed convergence field. In other words, the goal is to find the optimal mapping of the noisy convergence field to the noiseless convergence field. To this end, we apply two different machine learning (ML) approaches: generative adversarial networks (GANs) and diffusion models (DMs). These approaches are classified as deep generative models, which learn the probability distribution function of the target observable given the training data sets. In particular, the denoising process can be regarded as the image-to-image translation problem, where the input image is noisy observed convergence maps and the output image is noiseless true convergence maps. The deep generative models are effective at image-to-image translation problems and demonstrate high performance in practical tasks such as image inpainting, colourisation, and segmentation. Recently, these methods have also been applied to the noise reduction task in astrophysics and cosmology (e.g., Shirasaki2019; Moriwaki2021; Floss2024). Here, we overview the technical basics of GANs and DMs.
@@ 3.1.1 Generative adversarial network
First, we review the algorithm of GAN (Goodfellow2014). In particular, our model learns the mapping conditioned on the input convergence map, and thus, we utilise the conditioned version of GAN, i.e., conditional GAN [^1] In this work, we employ conditional GAN only and refer to conditional GAN as “GAN” for simplicity.. GANs consist of two competing networks: the generator $G$, which produces the fake images conditioned on the given image, and the discriminator $D$, which distinguishes fake and true images. At the same time, the discriminator compares the input image with the target image to determine whether the target image is a real image or a fake image generated by the generator. Through training with data sets, the model parameters $\theta$ are determined by optimising the Jensen–Shannon divergence and regularisation term:
@@
where $\mathbb{E}_{\bm{p}}$ denotes the averaging operation over the variable $\bm{p}$, $\bm{x}$ and $\bm{y}$ are the input and output data vector, respectively, and $\bm{z}$ is a random noise added to the lowest layer of the generator. Although it is possible to learn the transformation from the input to the output without noise $\bm{z}$, Gaussian noise $\bm{z}$ is added to prevent overfitting to the training data distribution and induce a diversity of output data. During training, gradient descent is applied alternately to optimise both $D$ and $G$ in each step. As described in the original implementation of GANs (Goodfellow2014), during training of $G$, instead of minimizing $\log(1-D(\bm{x},G(\bm{x},\bm{z})))$, we minimize $\log D(\bm{x},G(\bm{x},\bm{z}))$. Furthermore, to slow down the learning rate of $D$, the loss function is divided by two during the optimisation of $D$. In summary, we alternately minimize the loss functions $L_{G}$ and $L_{D}$, which are defined as
@@
In this work, we use pix2pix (Isola2017) implemented with PyTorch (PyTorch)  [^2] https://github.com/junyanz/pytorch-CycleGAN-and-pix2pix, which employs GAN to take two-dimensional images as data vectors. In addition to the loss function originally proposed in Isola2017 (Eqs. 10 and 11), we also train the model with alternative loss functions: LSGAN (Mao2016) and WGAN-gp (Gulrajani2017). These loss functions are proposed to improve the stability of training for generic GANs. However, we find that the model trained with the original loss function yields the best results. LSGAN stabilizes the training but the denoising results are worse than the original loss fuction. WGAN-gp leads to unstable training by breaking the balance between generator and discriminator losses, and the generated images are collapsed. Therefore, we employ the original loss function as fiducial choice. The details of the architectures of the networks are described in Appendix A.1.
While GANs can generate high-quality images quickly, they suffer from various issues due to the model structure. In general, the training of GANs is unstable due to the mode collapse and the vanishing gradient problem compared with other existing generative models such as variational auto-encoders (Kingma2014b). Furthermore, the generated images are less diverse because the generator tends to ignore the noise term and generate images only from the information of the conditioned images as training proceeds. Due to this problem, pix2pix does not provide the noise term in the input data vectors. Instead, the stochasticity appears only in dropouts of the generator network. However, the output images are still less diverse even with the treatment than other generative models, that is, GANs fail to learn the probability distribution of the target images.
@@ 3.1.2 Diffusion model
DM (Sohl-Dickstein2015; Ho2020; Song2021) has been proposed in order to circumvent the problems of GANs and enhance the quality of generated data. It is known that training of DM is more stable and can generate more diverse outputs from a single input data.
@@
where $\mathcal{N}(\bm{\mu},\bm{\Sigma})$ is the normal distribution with the mean $\bm{\mu}$ and the covariance $\bm{\Sigma}$, and $\bm{I}$ is the identity matrix. The probability distribution function $q(\bm{y}_{t}|\bm{y}_{t-1})$ of the noise-added data $\bm{y}_{t}$ given the data at the previous step $\bm{y}_{t-1}$ is expressed as
@@
where the mean $\bm{\mu}_{\theta}$ and the variance $\bm{\Sigma}_{\theta}$ are expressed as the deep neural networks with model parameters $\theta$. Then, let us consider the consistency of the joint probability distribution function of the series of the data $(\bm{y}_{1},\ldots,\bm{y}_{T})$. The joint probability distributions are given with forward process or reverse process:
$q(\bm{y}_{1},\ldots,\bm{y}_{T})$ $=q(\bm{y}_{0})\prod_{t=1}^{T}q(\bm{y}_{t}|\bm{y}_{t-1}),$ (19)
$q_{\theta}(\bm{y}_{1},\ldots,\bm{y}_{T})$ $=p(\bm{y}_{T})\prod_{t=1}^{T}q_{\theta}(\bm{y}_{t-1}|\bm{y}_{t}),$ (20)
where $p(\bm{y}_{T})$ is the prior of the final data and taken as unit normal distribution. The objective is to minimise the Kullback–Leibler divergence of the two distributions: $q(\bm{y}_{1},\ldots,\bm{y}_{T})$ and $q_{\theta}(\bm{y}_{1},\ldots,\bm{y}_{T})$.
In practice, for better convergence of the training process and sampling efficiency, Ho2020 proposed the simplified scheme; the objective function for optimisation is the reweighted variational lower-bound and the covariance is diagonal $\bm{\Sigma}_{\theta}=\sigma_{t}^{2}\bm{I}$. Finally, the loss function $L_{t}$ for the step $t$ is given as
@@
In this work, we employ Palette (Saharia2021) model, which implements DM to take images as inputs and outputs with PyTorch (PyTorch). Similarly to GANs, Palette takes the input image $\bm{x}$ used as the conditioned image, and thus, the model should be referred to as “conditional” DM. Conditional DM can be implemented by adding the input image as the argument of the noise vector, i.e.,
@@
DM can generate more high-quality and diverse data compared with GAN. Furthermore, the training process of DM is more stable because of the simple objective function (Eq. 21) in contrast to competing loss functions of the generator and discriminator in GAN. The drawback of DM is the computational cost of data generation. In general, the number of total diffusion steps $T$ is required to be $T\gtrsim\mathcal{O}(10^{3})$, which takes considerably longer than GANs.